OpenStudy (anonymous):

The area of circle is 32pi. what's the circumference

OpenStudy (barrycarter):

Hint: if it's are is 32pi, can you find the radius?

OpenStudy (unklerhaukus):

\[A_{\text{circle}}=\pi r^2\] \[C=2\pi r\]

OpenStudy (anonymous):

r^2=32 is it right

OpenStudy (anonymous):

how to get the r

OpenStudy (unklerhaukus):

right , now take the square root to get r

OpenStudy (anonymous):

r=sqrt 32

OpenStudy (anonymous):

4 sqrt 2

OpenStudy (unklerhaukus):

yep, now you can plug this into the formula for the circumference

OpenStudy (anonymous):

35.5 units

OpenStudy (anonymous):

Thank you so much for helping me

OpenStudy (unklerhaukus):

you've got the right answer

OpenStudy (unklerhaukus):

\[C=2\pi\times 4\sqrt 2\]\[=8\sqrt2\pi\]\[\approx 35.54\]

OpenStudy (anonymous):

Thanks